Write the molecular, ionic and net ionic equations for the reaction that occurs when aqueous solutions of K3PO4 and Sr(NO3)2.
1
Expert's answer
2012-07-03T08:44:34-0400
2 K3PO4 (aq) + 3 Sr(NO3)2 (aq) = Sr3(PO4)2 (s) + 6 KNO3 (aq)
6 K(+) + 2
PO4(3-) + 3 Sr(2+) + 6 NO3(-) = Sr3(PO4)2 + 6 K(+) + 6 NO3(-)
2 PO4(3-) + 3
Sr(2+) = Sr3(PO4)2